OpenStudy (anonymous):

A line has a slope of –3 and a y-intercept of (0, –1). What is the equation of the line that is parallel to the given line and passes through the point (–3, 1)? A.y = –3x – 8 B.y = 3x + 10 C.y = –3x + 4 D.y = 1/3x+2

OpenStudy (mathstudent55):

Start with the point-slope form of the equation of a line y = mx + b

OpenStudy (anonymous):

Line is parallel...therefore slope is same hence answer has to be from A or C! Check the equation which satisfies the point (-3,1) I guess it's A

OpenStudy (mathstudent55):

You are told that the given line has a slope of -3, and you are looking for a parallel line. Parallel lines have equal slopes. That means your line has a slope of -3 too.

OpenStudy (mathstudent55):

Now you can put the slope in the equation: y = mx + b y = -3x + b Now to see which equation is correct, replace x and y in the equation above by the coordinates of the given point (-3, 1). 1 = -3(-3) + b 1 = 9 + b -8 = b b = -8 Since you now know b, you can finally write the complete equation of the new line: y = -3x - 8

OpenStudy (anonymous):

thanks for the help :)

OpenStudy (mathstudent55):

You're welcome.